x 2+ 6xy+y 2 = 10 to prove that d 2y /dx 2 = 80/ (3x+y) 3

x2+6xy+y2  = 10  to prove that  d2y /dx2  =  80/  (3x+y)3

Grade:12

1 Answers

Khimraj
6 years ago
X2+6xy+y2 = 10
2x + 6y + 6x(dy/dx) + 2y(dy/dx) =0
(dy/dx) = -(x+3y)/(3x+y)
d2y/dx2 = -[(3x+y)(1+3dy/dx) – (x+3y)(3+dy/dx)]/(3x+y)2
put dy/dx = -(x+3y)/(3x+y)
d2y/dx2 = 80/  (3x+y)3
Hope it clears.